OpenStudy (anonymous):

Write -4x-2y=-6 in slope incept form.

hero (hero):

-4x - 2y = -6 -4x + 6 = 2y -2x +3 = y

OpenStudy (anonymous):

y=-2x+3

OpenStudy (anonymous):

why doesn't the -2 change to a positive 2

hero (hero):

Because I added 6 to both sides, then added 2y to both sides. I left the -4x. So when I divided both sides by 2, -4/2 = -2

OpenStudy (anonymous):

cool